OpenStudy (anonymous):

(1)/(w)=(2)/(w)-(1)/(3) solve

OpenStudy (anonymous):

1/w = 2/w - 1/3 1/w = 6-w / 3w 3w = 6w - w^2 w^2 - 3w = 0 w = 3 or 0